# GAS: Generative Auto-bidding with Post-training Search

### v1

@@ Abstract.
Auto-bidding is essential in facilitating online advertising by automatically placing bids on behalf of advertisers. Generative auto-bidding, which generates bids based on an adjustable condition using models like transformers and diffusers, has recently emerged as a new trend due to its potential to learn optimal strategies directly from data and adjust flexibly to preferences. However, generative models suffer from low-quality data leading to a mismatch between condition, return to go, and true action value, especially in long sequential decision-making. Besides, the majority preference in the dataset may hinder models’ generalization ability on minority advertisers’ preferences. While it is possible to collect high-quality data and retrain multiple models for different preferences, the high cost makes it unaffordable, hindering the advancement of auto-bidding into the era of large foundation models. To address this, we propose a flexible and practical Generative Auto-bidding scheme using post-training Search, termed GAS, to refine a base policy model’s output and adapt to various preferences. We use weak-to-strong search alignment by training small critics for different preferences and an MCTS-inspired search to refine the model’s output. Specifically, a novel voting mechanism with transformer-based critics trained with policy indications could enhance search alignment performance. Additionally, utilizing the search, we provide a fine-tuning method for high-frequency preference scenarios considering computational efficiency. Extensive experiments conducted on the real-world dataset and online A/B test on the Kuaishou advertising platform demonstrate the effectiveness of GAS, achieving significant improvements, e.g., 1.554% increment of target cost.
@@
copyright: acmlicensedjournalyear: 2025doi: xx.xxconference: the ACM Web Conference 2025; Auto-bidding; 2024isbn: 978-1-4503-XXXX-X/18/06submissionid: 123-A56-BU3ccs: Information systems Computational advertising
@@ 1. Introduction
The rapid digitalization of commerce has significantly expanded the reach and importance of online advertising platforms, making them essential for businesses to engage target audiences and boost sales (Wang and Yuan, 2015; Evans, 2009). With the vast number of impression opportunities, manually adjusting bid prices to optimize costs under budget and KPI constraints is impractical. To address this, ad platforms now offer auto-bidding services that automate the bidding process using advanced strategies (Balseiro et al., 2021; Deng et al., 2021; Ou et al., 2023). These strategies consider factors from immediate or historical bidding information, such as the distribution of impression opportunities and remaining budgets (Li and Tang, 2022). Besides, according to different types of advertisers, the strategy should take their different preferences into consideration. For example, brand advertisers, who aim for long-term growth and awareness, typically aim to show their ads to as many people as possible under constraints like average cost per impression. However, performance advertisers, who focus on maximizing the value of winning impressions, aim to maximize conversions with cost-per-conversion constraints (He et al., 2021). To meet these diverse needs, advertising platforms like Google, Facebook, and Alibaba provide tailored multiple bidding strategies for their customers (Google, 2021; Facebook, 2021; Alibaba, 2021). Additionally, in the face of the dynamic ever-changing advertising environment, strategies must be regularly optimized to stay closely aligned with customers’ preferences, thereby helping them achieve long-term commercial benefits (Guo et al., 2024).
@@
Reinforcement learning (RL) has long been a leading method for enhancing auto-bidding strategies by training agents with an advertising simulator or offline advertising logs. However, RL approaches are predominantly based on the Markovian decision process (MDP), which assumes that future states are determined solely by the current single step’s state and action. Recent statistical analyses (Guo et al., 2024) question this assumption in auto-bidding, revealing a strong correlation between the lengths of historical state sequences and subsequent states. This finding suggests that relying solely on the most recent state can lead to instability in the unpredictable online advertising environment. Moreover, the preference for the RL strategy is not easy to control. USCB (He et al., 2021) proposes calculating an optimal solution under multiple constraints using historical data and then training RL to refine the strategy. However, once deployed, the strategy’s preference is fixed, limiting interactivity and controllability. Therefore, this has led to a new trend in developing generative auto-bidding methods based on conditional generative models like transformers and diffusers. With a vector condition representing the preferences, these methods directly output an action or even a trajectory without MDP assumption. For example, decision transformers (Chen et al., 2021) can utilize extensive historical information for decision-making. Diffusers (Ajay et al., 2023) can directly generate a trajectory for planning given a condition. More importantly, generative models provide flexibility in controlling preferences by simply modifying their condition’s value during deployment. As large foundational models based on generative models have shown remarkable progress in applications like natural language processing and computer vision, exemplified by ChatGPT (OpenAI, 2024) and Stable Diffusion (Rombach et al., 2022), it is foreseeable that the auto-bidding field will also advance towards the foundation model era. This will involve developing a decision foundation model to learn optimal decision strategies directly from large-scale datasets.
@@ 4.1. Setup
Datasets. Different from the previous dataset preparation procedure that collects private bidding logs from a non-open-source advertising auto-bidding system, we utilize a new public large-scale real-world bidding dataset, termed AIGB, released by Alibaba (Xu et al., [n. d.]). This dataset, to our knowledge the largest available, includes over 2 million trajectories and offers a sparse version for more challenging cases. More details are in Appendix A.
@@
Baselines. We compare our method to a wide range of baselines involving both RL and generative methods: USCB (He et al., 2021), an online RL approach that adjusts parameters dynamically to the optimal bids; BCQ (Fujimoto et al., 2019), a typical offline RL method that updates policies with only access to a fixed dataset; CQL (Kumar et al., 2020), an offline RL method for learning conservative value function by regularizing Q-values; IQL (Kostrikov et al., 2022), an offline RL method that enables multi-step dynamic programming updates without querying out-of-sample actions; DiffBid (Guo et al., 2024), a generative method based on diffusers to generate bidding trajectories given conditions; DT (Chen et al., 2021), a sequential decision-making generative method based on the transformer; CDT (Liu et al., 2023), a DT-based method that considers a vector of multiple constraints; DT-score: a DT-based method that utilizes the reward function considering both the winning value and the KPI constraints.
Implementation Details. The hyperparameters of baselines are set based on the default values referenced in original papers, with further tuning performed to optimize performance. For GAS, it has two components: a base policy model and multiple QT networks. The base policy model could be any model and we choose the DT-score with hyperparameters referenced from the official code provided from (Xu et al., [n. d.]), which a smaller learning rate at $1e^{-5}$ for fine-tune specifically. For QT network, we employ 6 attention layers with 8 heads for each and set the hidden size as 512 for all layers with 14M parameters in total, which is lightweight. The total training steps are 400k. We adopt the AdamW optimizer (Loshchilov and Hutter, 2019) with a learning rate of $1e^{-4}$, and set the batch size as 128. During training, we utilize the PyTorch framework on two NVIDIA H100 GPUs. For more detailed hyperparameters of QT networks, please refer to Table 6 from Appendix B .
@@ 4.2. Performance Comparison with Baselines
In this experiment, we conducted a performance comparison of various baseline methods under different settings, including varying datasets and budget constraints in the MCB bidding. The dataset used for this evaluation is AIGB-2M and its sparse version, AIGB-sparse, and the budgets are set at 50%, 75%, 100%, 125%, and 150% of the maximum allowable budget. The results are indicated by the score as a comprehensive assessment of the performance. Our method is based on the baseline policy model DT-score.
@@ 4.3. Preference Alignment Performance
AIGB-2M Value $\uparrow$ 373 391 339 402 375 370 359
@@
AIGB-sparse Value $\uparrow$ 36.8 38.9 36.9 39.2 37.8 36.4 36.3
@@ 4.4. Ablation Study
Range of Search. The search range directly impacts the balance between exploration (trying out new actions to discover their potential) and exploitation (focusing on actions known to yield good results). A smaller search range means the algorithm will explore fewer actions, potentially missing out on some good actions but focusing more on refining the evaluation of the explored actions. A larger search range means the algorithm will explore more actions, increasing the chances of finding the best possible actions but at the cost of spending less time on each action, which might lead to less accurate evaluations. An extreme case could be to search the whole action space, while it could be low-efficiency. Considering the computational efficiency, we randomly search 5 actions within a search range and the ablation study results for different search ranges’ results are shown in Fig. LABEL:fig:range_of_search. The experimental results suggest that for the AIGB-2M algorithm, a 10% search range is better, providing the highest performance. Increasing the search range beyond 10% leads to diminishing returns, and the min-max approach is the least effective due to the limited action budget. More importantly, as the optimal performance is achieved with a small search range of around 10%, it indicates that a small refinement based on the base policy model could be good enough, leaving more space for a more advanced value assessment method.
@@
Effectiveness of QT. As the rollout process is skipped and approximated by the Q-value function, the accuracy of the Q-value function in predicting the expected return with the underlying policy becomes essential. Without information from the historical trajectory, predicting the rollout process given a state $s_{t}$ and action $a_{t}$ is unclear and highly random. This insight is supported by our experimental results in Table 3, where the performance of QT with historical trajectory as policy representation achieves much higher performance than the Q-value function based on a vanilla state-action pair.
We also compare QT to various methods without a Q-value net to testify to the effectiveness of MCTS search, including greedy search, and random/mean selection. Greedy search selects the top action with the highest predicted immediate reward, random selection randomly selects an action, and mean selection takes the average value of 5 random actions. As shown in Table 3, our method significantly outperforms other methods.
Table 3. Effectiveness of QT. Except for GAS-infer, other methods do not employ a QT during the search.
AIGB-2M DT-score Greedy Mean Random GAS w/o QT GAS-infer
@@ 5. Live Experiment
Our online A/B testing is conducted for five full days. For each MCB campaigns, 25% budget and traffic are allocated to the baseline bidding model and GAS. The results are shown in Table 4. The experiment results show that GAS improved impressions by +0.934%, cost by +0.953%, target cost by 1.554%, and overall ROI by +0.595%. All metrics showed significant improvements.
@@ 6. Related Work
Auto-bidding and Offline Reinforcement Learning. To eliminate the need for manual intervention, auto-bidding autonomously manages ad placement bids on behalf of advertisers to optimize key performance indicators (KPIs) like clicks. Initially, classic control methods like PID (Chen et al., 2011) were used to optimize budget spending using predefined curves, while OnlineLP (Yu et al., 2017) adjusted bids based on traffic predictions. With increasing complexity of the online bidding environments, RL algorithms such as USCB (He et al., 2021), SORL (Mou et al., 2022), and MAAB (Wen et al., 2022) became crucial for decision-making. Especially offline RL methods, which learn policies from existing datasets without online interaction, have shown significant success. Notable methods include BCQ (Fujimoto et al., 2019), which narrows the action space to guide the agent towards on-policy behavior; CQL (Kumar et al., 2020), which regularizes Q-values for conservative estimates; and IQL (Kostrikov et al., 2022), which enables multi-step dynamic programming updates without querying Q values of out-of-sample actions during training. However, these methods are constrained by the MDP assumption, whereas generative models show greater potential for auto-bidding.
@@
Generative Models. Generative models seek to learn the underlying distribution of a given dataset or establish conditional probability distributions between variables. Deep generative models like VAEs (Doersch, 2016), Flows (Pan et al., 2023), and GANs (Goodfellow et al., 2020) have ushered in a new era of mastering complex data distributions by representing high-dimensional data in a controllable Gaussian form. Recently, the decision transformer (DT) (Chen et al., 2021) has been employed to model complex decision-making distribution as an auto-regressive model based on the transformer architecture. Diffusion models (Ho et al., 2020) generate a high-quality sample based on a condition by iterative denoising via a reverse process. For auto-bidding, DiffBid (Guo et al., 2024) employs a conditional diffusion model to generate a long trajectory based on a condition like return and then uses an inverse dynamic model to learn how to map the current state to the predicted next state.
@@
Preference Alignment. There are two major ways to finetune a base model to specific preferences, including supervised fine-tuning (SFT) and preference optimization from human feedback. For SFT (Ouyang et al., 2022b), it constructs high-quality trajectories on specific domains to enhance the base model’s ability in these preferences, which is stable for finetuning. For the latter, it needs to collect a query dataset where typically two responses are manually annotated as chosen or rejected for a prompt. Reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022a), as a pioneering approach, uses a reward model and proximal policy optimization (PPO) (Schulman et al., 2017) to further refine the model. Direct preference optimization (DPO) (Rafailov et al., 2023) and its variants (Azar et al., 2024; Xu et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024; Park et al., 2024) then proposes to directly fine-tune a model using preference query data, bypassing the reward modeling phase and the RL optimization process. Recently, some post-training methods have been proposed to align the preference even without finetuning by using reward models only (Khanov et al., 2024; Li et al., 2024).
@@ 7. Conclusion
However, there are several limitations in this study. Firstly, the simplifications made in the MCTS process, such as approximating the expansion and simulation steps, may not fully capture the complexities of real-world bidding scenarios where other advertisers’ behaviors are unpredictable and the transition dynamic of the system is changing over time. Secondly, the fine-tune version of GAS is more computationally efficient than the inference version, but its performance is still limited, which needs more advanced and effective fine-tune methods.
Appendix A DETAILED DESCRIPTIONS OF DATASETS
Table 5. The Parameters of AIGB-2M and AIGB-sparse
Parameters Values
AIGB-2M AIGB-sparse
Number of trajectories 2089373 1718490

### v2

@@ Abstract.
Auto-bidding is essential in facilitating online advertising by automatically placing bids on behalf of advertisers. Generative auto-bidding, which generates bids based on an adjustable condition using models like transformers and diffusers, has recently emerged as a new trend due to its potential to learn optimal strategies directly from data and adjust flexibly to preferences. However, generative models suffer from low-quality data leading to a mismatch between the condition, like return to go, and true action value, especially in long sequential decision-making. Besides, the majority preference in the dataset may hinder models’ generalization ability on minority advertisers’ preferences. While it is possible to collect high-quality data and retrain multiple models for different preferences, the high cost makes it unaffordable, hindering the advancement of auto-bidding into the era of large foundation models. To address this, we propose a flexible and practical Generative Auto-bidding scheme using post-training Search, termed GAS, to refine a base policy model’s output and adapt to various preferences. We use weak-to-strong search alignment by training small critics for different preferences and an MCTS-inspired search to refine the model’s output. Specifically, a novel voting mechanism with transformer-based critics trained with policy indications could enhance search alignment performance. Additionally, utilizing the search, we provide a fine-tuning method for high-frequency preference scenarios considering computational efficiency. Extensive experiments conducted on the real-world dataset and online A/B test on the Kuaishou advertising platform demonstrate the effectiveness of GAS, achieving significant improvements, e.g., 4.60% increment of target cost. Code is available here.
@@
journalyear: 2025copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3715226isbn: 979-8-4007-1331-6/2025/04ccs: Information systems Computational advertising
@@ 1. Introduction
The rapid digitalization of commerce has expanded online advertising platforms’ reach, making them essential for engaging audiences and boosting sales (Wang and Yuan, 2015; Evans, 2009). Due to numerous impression opportunities, manually adjusting bids to optimize costs under budget and KPI constraints is impractical. To address this, ad platforms now offer auto-bidding services that automate the bidding process using advanced strategies (Balseiro et al., 2021; Deng et al., 2021; Ou et al., 2023). These strategies consider factors from immediate or historical bidding information, such as the distribution of impression opportunities and remaining budgets (Li and Tang, 2022). Besides, according to different types of advertisers, the strategy should take their different preferences into consideration. For example, brand advertisers, who aim for long-term growth and awareness, typically aim to show their ads to as many people as possible under constraints like average cost per impression. However, performance advertisers, who focus on maximizing the value of winning impressions, aim to maximize conversions with cost-per-conversion constraints (He et al., 2021). To meet these diverse needs, advertising platforms like Google, Facebook, and Alibaba provide tailored multiple bidding strategies for their customers (Google, 2021; Facebook, 2021; Alibaba, 2021). Additionally, facing the dynamic advertising environment, strategies must be regularly optimized to stay closely aligned with customers’ preferences, thereby achieving long-term commercial benefits (Guo et al., 2024).
@@
Reinforcement learning (Sutton, 2018; Pan et al., 2020) (RL) has long been a leading method for enhancing auto-bidding strategies by training agents with an advertising simulator or offline advertising logs. However, RL approaches are predominantly based on the Markovian decision process (MDP), which assumes that future states are determined solely by the current single step’s state and action. Recent statistical analyses (Guo et al., 2024) question this assumption in auto-bidding, revealing a strong correlation between the lengths of historical state sequences and subsequent states. This finding suggests that relying solely on the most recent state can lead to instability in the unpredictable online advertising environment. Moreover, the preference for the RL strategy is not easy to control. USCB (He et al., 2021) proposes calculating an optimal solution under multiple constraints using historical data and then training RL to refine the strategy. However, once deployed, the strategy’s preference is fixed, limiting interactivity and controllability. Therefore, this has led to a new trend in developing generative auto-bidding methods based on conditional generative models like transformers and diffusers. With a vector condition representing the preferences, these methods directly output an action or even a trajectory without MDP assumption. For example, decision transformers (Chen et al., 2021) can utilize extensive historical information for decision-making. Diffusers (Ajay et al., 2023) can directly generate a trajectory for planning given a condition. More importantly, generative models provide flexibility in controlling preferences by simply modifying their condition’s value during deployment. As large foundational models based on generative models have shown remarkable progress in applications like natural language processing and computer vision, exemplified by ChatGPT (OpenAI, 2024) and Stable Diffusion (Rombach et al., 2022), it is foreseeable that the auto-bidding field will also advance towards the foundation model era. This will involve developing a decision foundation model to learn optimal decision strategies directly from large-scale datasets.
@@ 4.1. Setup
Datasets. Different from the previous dataset preparation procedure that collects private bidding logs from a non-open-source advertising auto-bidding system, we utilize a new public large-scale real-world bidding dataset, termed AuctionNet, released by Alibaba (Su et al., 2024). This dataset, to our knowledge the largest available, includes over 500 million records and offers a sparse version for more challenging cases. More details are in Appendix A.
@@
Baselines. We compare our method to a wide range of baselines involving both RL and generative methods. For RL, we compare with online RL approach USCB (He et al., 2021), offline approaches BCQ (Fujimoto et al., 2019), CQL (Kumar et al., 2020), and IQL (Kostrikov et al., 2022). For generative methods, we compare with DiffBid (Guo et al., 2024), a generative method based on diffusers, and Decision Transformers DT (Chen et al., 2021) and its variants CDT (Liu et al., 2023), a DT-based method that considers a vector of multiple constraints, DT-score: a DT-based method that utilizes the reward function considering both the winning value and the KPI constraints.
Implementation Details. The hyperparameters of baselines are set based on the default values referenced in original papers, with further tuning performed to optimize performance. For GAS, it has two components: a base policy model and multiple QT networks. The base policy model could be any model and we choose the DT-score with hyperparameters referenced from the official code provided from (Xu et al., [n. d.]), which a smaller learning rate at $1e^{-5}$ for fine-tune specifically. For QT network, we employ 6 attention layers with 8 heads for each and set the hidden size as 512 for all layers with 14M parameters in total, which is lightweight. The total training steps are 400k. We adopt the AdamW optimizer (Loshchilov and Hutter, 2019) with a learning rate of $1e^{-4}$, and set the batch size as 128. During training, we utilize the PyTorch framework on two NVIDIA H100 GPUs. For more detailed hyperparameters of QT networks, please refer to Table 6 from Appendix A.2 .
@@ 4.2. Performance Comparison with Baselines
In this experiment, we conducted a performance comparison of various baseline methods under different settings, including varying datasets and budget constraints in the MCB bidding. The dataset used for this evaluation is AuctionNet and its sparse version, and the budgets are set at 50%, 75%, 100%, 125%, and 150% of the maximum allowable budget. The results are indicated by the score as a comprehensive assessment of the performance. Our method is based on the baseline policy model DT-score.
@@ 4.3. Preference Alignment Performance
AuctionNet Value $\uparrow$ 373 391 339 402 375 370 359
@@
AuctionNet-sparse Value $\uparrow$ 36.8 38.9 36.9 39.2 37.8 36.4 36.3
@@ 4.4. Ablation Study
Range of Search. The search range directly impacts the balance between exploration (trying out new actions to discover their potential) and exploitation (focusing on actions known to yield good results). A smaller search range means the algorithm will explore fewer actions, potentially missing out on some good actions but focusing more on refining the evaluation of the explored actions. A larger search range means the algorithm will explore more actions, increasing the chances of finding the best possible actions but at the cost of spending less time on each action, which might lead to less accurate evaluations. An extreme case could be to search the whole action space, while it could be low-efficiency. Considering the computational efficiency, we randomly search 5 actions within a search range and the ablation study results for different search ranges’ results are shown in Fig. LABEL:fig:range_of_search. The experimental results suggest that for the AuctionNet dataset, a 10% search range is better, providing the highest performance. Increasing the search range beyond 10% leads to diminishing returns, and the min-max approach is the least effective due to the limited action budget. More importantly, as the optimal performance is achieved with a small search range of around 10%, it indicates that a small refinement based on the base policy model could be good enough, leaving more space for a more advanced value assessment method.
@@
Effectiveness of QT. As the rollout process is approximated by the Q-value function, the accuracy of the Q-value function in predicting the expected return with the underlying policy becomes essential. Without historical trajectory information, predicting the rollout process given a state $s_{t}$ and action $a_{t}$ is unclear and highly random. This insight is supported by our experimental results in Table 3, where the performance of QT with historical trajectory as policy representation achieves much higher performance than the Q-value function based on a vanilla state-action pair. We also compare QT to various methods without a Q-value net to testify to the effectiveness of MCTS search, including greedy search, and random/mean selection. Greedy search selects the top action with the highest predicted immediate reward, random selection randomly selects an action, and mean selection takes the average value of 5 random actions. As shown in Table 3, our method significantly outperforms other methods.
Table 3. Effectiveness of QT.
AuctionNet DT-score Greedy Mean Random GAS w/o QT GAS-infer
@@ 5. Live Experiment
Our online A/B testing is conducted for six full days. For each MCB campaigns, 25% budget and traffic are allocated to the baseline bidding model and GAS. The results are shown in Table 4. The experiment results show that GAS improved impressions by +1.65%, cost by +0.94%, target cost by 4.60%, and overall ROI by +3.62%. All metrics showed significant improvements.
@@ 6. Related Work
Auto-bidding and Offline Reinforcement Learning. Initially, classic control methods like PID (Chen et al., 2011) were used to optimize budget spending using predefined curves, while OnlineLP (Yu et al., 2017) adjusted bids based on traffic predictions. With increasing complexity of the online bidding environments, RL algorithms such as USCB (He et al., 2021), SORL (Mou et al., 2022), and MAAB (Wen et al., 2022) became crucial for decision-making. Especially offline RL methods, which learn policies from existing datasets without online interaction, have shown significant success. Notable methods include BCQ (Fujimoto et al., 2019), which narrows the action space to guide the agent towards on-policy behavior; CQL (Kumar et al., 2020), which regularizes Q-values for conservative estimates; and IQL (Kostrikov et al., 2022), which enables multi-step dynamic programming updates without querying Q values of out-of-sample actions during training. However, these methods are constrained by the MDP assumption, whereas generative models show greater potential for auto-bidding.
@@
Generative Models. Generative models learn the data distribution of a given dataset or establish conditional probability distributions between variables. Deep generative models like VAEs (Doersch, 2016), Flows (Pan et al., 2023), and GANs (Goodfellow et al., 2020) have ushered in a new era of mastering complex data distributions by representing high-dimensional data in a controllable Gaussian form. Recently, the decision transformer (DT) (Chen et al., 2021; Liu et al., 2024; Gao et al., 2025) has been employed to model complex decision-making distribution as an auto-regressive model. Diffusion models (Ho et al., 2020) generate samples based on a condition by iterative denoising via a reverse process. For auto-bidding, DiffBid (Guo et al., 2024) employs a conditional diffusion model to generate a long trajectory based on a condition like return and then uses an inverse dynamic model to learn how to map the current state to the predicted next state.
@@
Preference Alignment. There are two major ways to finetune a base model to specific preferences, including supervised fine-tuning (SFT) and preference optimization from human feedback. For SFT (Ouyang et al., 2022b), it constructs high-quality trajectories on specific domains to align the base model to these preferences. For the latter, it needs to learn a preference model. Reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022a), as a pioneering approach, uses a preference model and proximal policy optimization (PPO) (Schulman et al., 2017) to further refine the model. Direct preference optimization (DPO) (Rafailov et al., 2023) and its variants (Azar et al., 2024; Xu et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024; Park et al., 2024) then proposes to directly fine-tune a model using preference query data, bypassing the reward modeling phase and the RL optimization process. Recently, some post-training methods have been proposed to align the preference even without finetuning by using reward models only (Khanov et al., 2024; Li et al., 2024).
@@ 7. Conclusion
However, there are several limitations in this study. Firstly, the simplifications made in the MCTS process, such as approximating the expansion and simulation steps, may not fully capture the full complexities of real-world bidding scenarios. Secondly, the fine-tune version of GAS is more computationally efficient than the inference version, but its performance is still limited, which needs more advanced and effective fine-tune methods.
Acknowledgements.
This research is supported by the National Research Foundation, Singapore, under its Competitive Research Programme (Grant No. NRF-CRP23-2019-0006). Additionally, Shuai Mao and Yunjian Xu are supported in part by the General Research Fund (GRF) project 14200720 of the Hong Kong University Grants Committee, and the National Natural Science Foundation of China (NSFC) Project 62073273.
Appendix A Appendix
A.1. Datset Details
The datasets consisting of AuctionNet and AuctionNet-sparse, where AuctionNet-sparse is a sparse version of AuctionNet with less conversions . Each dataset consists of 21 advertising delivery periods, with each period containing approximately 500000 impression opportunities, divided into 48 intervals. Detailed parameters are shown in Table 5. Each advertiser will bid for all impressions.
Table 5. The Parameters of AuctionNet and AuctionNet-sparse.
Params AuctionNet AuctionNet-Sparse
Trajectories 479,376 479,376
Delivery Periods 9,987 9,987